mathslover (mathslover):

The length and breadth of a thin rectangular sheet, measured using a metre scale, are 18.2 cm and 14.5 cm respectively. Find the uncertainty or error in the estimation of the area of the rectangular sheet.

mathslover (mathslover):

@UnkleRhaukus @hartnn

mathslover (mathslover):

@harsh314

OpenStudy (anonymous):

18.2 +/-0.1 or 14.5+/- 0.1 that's the least count

mathslover (mathslover):

Yes but my small (may be silly) , why 0.1 there?

hartnn (hartnn):

because the least count of meter scale is 0.1 cm

mathslover (mathslover):

Yes I got it now... thanks @hartnn

mathslover (mathslover):

So it will be : \[\large{\frac{0.1}{18.2}\pm \frac{0.1}{14.5}}\]

mathslover (mathslover):

as : \[\large{\delta Z = \frac{\delta A}{A} +\frac{\delta B}{\B}}\]

mathslover (mathslover):

\[\large{\delta Z = \frac{\delta A}{A} +\frac{\delta B}{B}}\]

mathslover (mathslover):

correct ? As we were talking about multiplication so the applied formula will be mentioned above

hartnn (hartnn):

not sure what u wrote is correct(dimensions not matching).....which reference did u use ?

mathslover (mathslover):

Let me check wait.